Periodically automobile firms offer incentives to spur sales.  The current one provided by GM, which apply to a $12,000 loan you are considering, provides for 2.9% financing with a 36-month loan.  Alternatively, a rebate of $1000 is available.  If you take the rebate you would have to finance the remaining $11,000 at an interest rate of 10% for 36 months.  Which incentive is worth more to you?  (GM financing)
Answer #1

Answer:

As both incentive requires monthly payment for 36 months, we can calculate monthly payment required for both of them and compare. The incentive with lower monthly payment will be worth more to you.

Option: GM Financing

Loan amount = $12,000

Monthly Interest = 2.9% / 12

Number of months = 36

Monthly Payment = PMT (rate, nper, pv, fv, type) = PMT (2.9%/12, 36, -12000, 0, 0) = $348.45

Option: Rebate

Loan amount = $12,000 - 1000 = $11,000

Monthly Interest = 10% / 12

Number of months = 36

Monthly Payment = PMT (rate, nper, pv, fv, type) = PMT (10%/12, 36, -11000, 0, 0) = $354.94

As such monthly payment of 'GM Financing' is lower at $348.45 by (354.94 - 348.45 =) $6.49 per month.

GM financing is worth more.